OpenStudy (anonymous):

what is the slope of the line that passes through the given points? (3,2) and (5,12) 1) 5 2) -5 3) -1/5 4) 1/5 i belive the answer is 1/5

OpenStudy (texaschic101):

let me see if you are correct... slope = (y2 - y1) / (x2 - x1) (3,2)...x1 = 3 and y1 = 2 (5,12)...x2 = 5 and y2 = 12 now sub slope = (12 - 2) / (5 - 3) slope = 10/2 slope = 5 nope...the slope is 5 not 1/5

OpenStudy (texaschic101):

do you understand how I got that ?

OpenStudy (anonymous):

ohhh i forgot to subtract i believe, silly mistakes, but thankyou so much! @texaschic101

OpenStudy (texaschic101):

no problem....don't feel bad...if I had $1 for every mistake I made, I would be rich...lol